You bought a parallel plate capacitor for the flash of your developed camera. It consists of two rectangular plates, each with area of 4.5 cm2, that are separated by a 2.00-mm thick ordinary glass. You then connected it with a 12.0 V battery. How much energy is stored in the capacitor?


1
Expert's answer
2022-02-18T08:22:11-0500

Given:

"d=0.002\:\rm m"

"A=4.5*10^{-4}\:\rm m^2"

"V=12\:\rm V"

"k=5"

The capacitance


"C=\frac{k\epsilon_0A}{d}=\frac{5*8.85*10^{-12}*4.5*10^{-4}}{0.002}\\
=9.96*10^{-12}\:\rm F"


The energy of electric field

"W=\frac{CV^2}{2}=\frac{9.96*10^{-12}*12^2}{2}=7.2*10^{-10}\:\rm J"